OpenStudy (anonymous):

A recipe for fruit punch calls for 2 and 1 over 2 cups of orange juice,1 and 1 over 3 cups of pineapple, and 1 and 3 over 4 cups of soda water. If June makes 4 times the recipe,how many cups of fruit punch will she have?

OpenStudy (mattfeury):

Please try to just post replies once :). Thanks.

OpenStudy (anonymous):

haha! i know i just wanted someone to answer my question

OpenStudy (mattfeury):

Yea it just helps keep things clean and lets people help other questions too. I'm confused by the verbiage here though. What does "2 and 1 over 2 cups" mean? does that mean 2 1/2 ?

OpenStudy (anonymous):

yeah. just didn't know how to write it

OpenStudy (mattfeury):

ok so first we find out how much punch she has for the recipe. then we can just multiply that by 4.

OpenStudy (mattfeury):

\[2\frac{1}{2} + 1\frac{1}{3} + 1\frac{3}{4}\] you'll need to make those into improper fractions, then find a common denominator to add them.

OpenStudy (anonymous):

so i have to times the fraction?

OpenStudy (mattfeury):

the answer to the fraction addition will be another fraction. yes, then multiply that answer by 4

OpenStudy (mattfeury):

the answer to above is the amount of punch made by the recipe.

OpenStudy (anonymous):

i got it thanks for t he help